# Semantics-Oriented Multitask Learning for DeepFake Detection: A Joint Embedding Approach

### v1

@@ IV Proposed Method: SJEDD
In this section, we describe Semantics-oriented Joint Embedding DeepFake Detector (SJEDD) in detail, our joint embedding approach for semantics-oriented multitask DeepFake detection, including 1) general problem formulation, 2) joint embedding for detection, and 3) bi-level optimization for parameter estimation. The overall structure of SJEDD is shown in Fig. 3.
@@ IV-A Problem Formulation
Given an input face image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$, we assign to it a vector of binary labels $\bm{y}=\{0,1\}^{C}$, indicating one primary and $C-1$ auxiliary tasks. Specifically, $y_{1}=1$ signifies a forged face, and $y_{1}=0$ means the opposite. For $i>1$, $y_{i}=1$ indicates that the $i$-th face attribute/region is fake, and $y_{i}=0$ otherwise. Next, we define an unnormalized probability distribution over the label hierarchy [85]:
@@
where $s_{i}(\bm{x})$ represents the raw score of the detector for the $i$-th label, and $\mathbb{I}[\cdot]$ is an indicator function. Eq. (1) encodes two label relations: 1) if the $j$-th node is activated (i.e., $y_{j}=1$), at least one of its parent nodes, collectively indexed by $\mathcal{P}_{j}$, must also be activated; and 2) if the $i$-th node is activated, at least one of its child nodes, indexed by $\mathcal{C}_{i}$, must also be activated. We then normalize Eq. (1) to obtain the joint probability
@@ IV-B Joint Embedding
We first introduce text templates to encode the label hierarchy and then implement joint embedding using vision-language correspondence to estimate the raw scores for each label.
@@ IV-B1 Text Templates
For the root node face, the text template is defined as “a photo of a fake face.” At the global face attribute level, we use the template “A photo of a face with the global attribute of $\{g\}$ altered,” where $g$ belongs to the set of global face attributes {expression, identity, physical_inconsistency}. At the level of the local face region, the template is defined as “A photo of a face with the local $\{l\}$ region altered,” where $l$ is in the set of local face regions {eye, eyebrow, lip, mouth, nose, skin}.
@@ IV-B2 Vision-Language Correspondence
We construct two encoders: an image encoder $\bm{f}_{\bm{\phi}}:\mathbb{R}^{H\times W\times 3}\mapsto\mathbb{R}^{N\times 1}$, parameterized by $\bm{\phi}$, which computes the visual embedding $\bm{f}_{\bm{\phi}}(\bm{x})$ for an input face image $\bm{x}$; and a text encoder $\bm{g}_{\bm{\varphi}}:\mathcal{T}\mapsto\mathbb{R}^{N\times 1}$, parameterized by $\bm{\varphi}$, which computes all textual embeddings $\{\bm{g}_{\bm{\varphi}}(\bm{t}_{i})\}_{i=1}^{C}$. Here, $\bm{t}_{i}$ denotes the text template for the $i$-th label, and $\mathcal{T}$ encompasses all candidate text templates. We estimate the raw score in Eq. (1) by
@@ IV-C Bi-level Optimization
Our goal is to optimize the joint embedding function, parameterized by $\bm{\phi}$ and $\bm{\varphi}$, which outputs the raw scores for all labels, based on which their marginal probabilities $\hat{p}(y_{i}|\bm{x};\bm{\phi},\bm{\varphi})$ can be computed. Given a training minibatch $\mathcal{B}_{\mathrm{tr}}=\{\bm{x}^{(k)},\bm{y}^{(k)},\mathcal{Z}^{(k)}\}_{k=1}^{K}$, where $\bm{y}^{(k)}=\{0,1\}^{C}$ is the complete ground-truth vector and $\mathcal{Z}^{(k)}\subseteq\{1,\ldots,C\}$ is the index set of observed labels, we minimize an overall loss function that is a linear weighted summation of the individual losses for each observed label:
@@
where $\bm{\lambda}=[\lambda_{1},\ldots,\lambda_{C}]^{\intercal}$ is the loss weighting vector to be optimized together with the model parameters $\bm{\phi}$ and $\bm{\varphi}$.
While the de facto cross-entropy function can be used to implement $\ell(\cdot)$ in Eq. (5) for parameter estimation, it is unbounded from above [44]. This can result in excessive penalties for challenging training examples, hindering model capacity allocation and loss weighting adjustment in multitask learning. To address this issue, we resort to the fidelity loss [44], derived from quantum physics, to measure the statistical distance between two states of a quantum:
@@
To automatically adjust the loss weighting vector $\bm{\lambda}$ and prioritize the primary DeepFake detection task, we employ bi-level optimization [43]. At the outer level, we minimize the primary task concerning $\bm{\lambda}$ on the validation minibatch $\mathcal{B}_{\mathrm{val}}$ to measure and promote generalization. At the inner level, we minimize the overall loss concerning the model parameters $\bm{\phi}$ and $\bm{\varphi}$ on the training minibatch $\mathcal{B}_{\mathrm{tr}}$. This gives rise to the following bi-level optimization problem:
@@
where ||$\cdot$|| ⋅ | denotes the cardinality of a set. As suggested in [41], we share a common loss weighting for tasks within the same level. Specifically, we learn to adjust three $\lambda_{i}$’s: one for the primary task, one for tasks at the global image attribute level, and one for the local image region level. We sample training and validation data as different minibatches in the same training dataset. We utilize the finite difference method as an efficient approximation to solve Problem (7).
V Experiments
In this section, we first describe the experimental setups and then thoroughly compare the proposed SJEDD with contemporary DeepFake detectors. In addition, we perform a series of ablation experiments to validate the efficacy of SJEDD from both the model and the data perspectives.
V-A Experimental Setups
V-A1 Datasets
We adopt the widely used FF+ [18] and newly established FFSC [41] for training. FF++ contains $1,000$ real videos, manipulated by four DeepFake methods: Deepfakes (DF) [1], Face2Face (F2F) [84], FaceSwap (FS) [83], and NeuralTexures (NT) [4]. These videos come with three compression levels, and our experiments are conducted at the slight compression level, as recommended by [24, 66, 28]. We expand FF++ using the proposed dataset expansion technique described in Sec. III. In contrast, FFSC [41] organizes twelve face manipulations into a semantic hierarchical graph with five global face attribute nodes and six local face region nodes and requires no dataset expansion.
We test DeepFake detectors on four independent DeepFake datasets: CDF [46], FaceShifter (FSh) [45], DeeperForensics-1.0 (DF-1.0) [47], and DeepFake Detection Challenge (DFDC) [48]. Notably, CDF and DFDC are particularly challenging. CDF is difficult due to its high-quality, realistic face forgeries. DFDC poses challenges with fake videos created from real-world footage, often featuring faces in low lighting, with extreme poses, and under occlusions.
V-A2 Implementation Details
SJEDD relies on the CLIP model [42] for joint embedding, where the image encoder is ViT-B/32 [86] and the text encoder is GPT-2 [87]. SJEDD is fine-tuned by minimizing the Problem (7) using AdamW [88] with a decoupled weight decay of $31e-$ 3 and a minibatch size of $32$. The initial learning rate is set to $76e-$ 7, following a cosine annealing schedule [89]. The loss weightings are all initialized to one and optimized using Adam [90] with a learning rate of $31e-$ 3. Throughout training, the temperature parameter $\tau$ in Eq. (4) is fixed at $10$. Following the practices in [25, 19, 27], we randomly introduce one of five real perturbations to each input image with a probability of $0.3$: patch substitution, additive white Gaussian noise, Gaussian blurring, pixelation, and JPEG compression. These perturbations are constrained to be mild to preserve face semantics (i.e., label-preserving). The training process runs for a maximum of $20$ epochs, taking approximately $21$ hours on a single NVIDIA RTX 3090 GPU.
TABLE II: AUC results of the proposed SJEDD against eighteen DeepFake detectors in the cross-dataset setting. All models are trained using the training set of FF++ (and its augmented versions). We retrain CNND, Face X-ray, MADD, FRDM, SLADD, DCL, and SO-ViT-B to match the performance reported in their original papers. For PCL, SFDG, VLFFD, and F2Trans-B, we directly use the results from their corresponding papers due to the lack of publicly available implementations. The last column shows the mean AUC numbers, including/excluding the FF++ test set. The best results are highlighted in bold
@@ V-A3 Competing Methods
We compare SJEDD against eighteen state-of-the-art DeepFake detectors: CNND [19], Face X-ray [21], PCL [22], MADD [27], Lip-Forensics [28], FRDM [29], RECCE [23], SBI [30], ICT [31], SLADD [24], DCL [32], CADDM [25], SFDG [34], VLFFD [33], F2Trans-B [35], MRL [36], CFM [37], and SO-ViT-B [41]. CNND serves as a typical baseline model. MADD employs attention mechanisms for feature selection and aggregation. FRDM and F2Trans-B expose face forgeries through high-frequency analysis. DCL and CFM leverage contrastive learning, while SFDG and MRL utilize graph learning to improve DeepFake detection. Face X-ray, PCL, SLADD, and CADDM incorporate manipulation localization as a manipulation-oriented auxiliary task to facilitate DeepFake detection. SLADD and VLFFD also include $C$-way manipulation classification. RECCE suggests reconstructing face images additionally to learn generative face features. Various empirical losses such as region overlap loss [27], additive margin softmax loss [29], InfoNCE loss [32], and ArcFace loss [31] have also been applied to enhance discriminative feature learning. Furthermore, Face X-ray, PCL, SBI, ICT, SLADD, and VLFFD exploit different data augmentation strategies to enhance detection performance, including face swapping between real faces [21, 22, 30, 31] and manipulation of the local face region [24, 33]. Lip-Forensics and ICT rely on high-level lipreading and face identity features, respectively. Like the proposed SJEDD, SO-ViT-B employs a semantics-oriented multitask learning paradigm but differs in both data and model aspects.
@@ V-A4 Evaluation Metrics
In line with [28, 29, 30, 31, 24, 25, 23, 32, 27, 26], we evaluate the performance of a DeepFake detector using the area under the receiver operating characteristic curve (AUC (%)).
@@ V-B1 Training on FF++
Cross-dataset evaluation. Table II shows the AUC results, revealing several interesting observations. First, many detectors struggle with generalization in the cross-dataset setting. In contrast, SJEDD performs consistently well on all five DeepFake datasets, particularly excelling in challenging datasets such as CDF, DFDC, and FFSC, and achieves the highest mean AUC results. Second, all methods show saturated performance on FF++. Third, while SBI reports a high AUC of $93.18\%$ on CDF, it performs poorly on DF-1.0 and DFDC. Similarly, PCL achieves an exceedingly high AUC of $99.40\%$ on DF-1.0 but generalizes poorly on DFDC. This performance discrepancy may be due to heavily relying on manipulation-specific features, such as statistical inconsistencies in color appearances and landmark locations. Fourth, all competing methods score relatively low on DFDC, likely resulting from the significant domain shift caused by different imaging conditions and manipulation techniques. SJEDD achieves an impressive AUC of $84.14\%$ in DFDC, demonstrating the superior transferability of features learned through our semantics-oriented joint embedding paradigm. Last, adhering to the same semantics-oriented design philosophy, SJEDD outperforms its predecessor, SO-ViT-B, confirming the changes in model design (i.e., from discriminative learning to joint embedding) and loss function (i.e., from cross-entropy loss to fidelity loss in bi-level optimization).
TABLE III: AUC results in the cross-manipulation test setting. DF, F2F, FS, and NT represent Deepfakes [1], Face2Face [84], FaceSwap [83], and NeuralTexures [4], respectively. Gray numbers represent within-manipulation results
@@
SJEDD (Ours) 99.88 95.14 97.31 98.51 97.46
Cross-manipulation evaluation. We perform the cross-manipulation evaluation of SJEDD against six representative DeepFake detectors: MADD [27], RECCE [23], DCL [32], VLFFD [33], CFM [37], and SO-ViT-B [41]. Specifically, we train on one manipulation type from FF++ [18] (including additional augmented data) and test on the remaining manipulations from FF++. As shown in Table III, SJEDD surpasses other methods by clear margins, which we believe arises from two design choices. First, our semantics-oriented formulation helps SJEDD learn generalizable features across different manipulations that alter the same face attributes (e.g., NT $\rightarrow$ F2F, both modifying the global expression attribute and the local mouth and lip regions). Second, SJEDD models the joint probability distribution of face attributes (see Eq. (1)), thus promoting generalization across different face attributes (e.g., DF: identity $\rightarrow$ NT: expression). Once again, SJEDD outperforms its predecessor SO-ViT-B, highlighting the effectiveness of the joint embedding formulation and the fidelity loss in bi-level optimization.
Cross-attribute evaluation. As suggested in FFSC [41], we adopt two semantics-oriented testing protocols: Protocol-1, which focuses on generalization to novel manipulations for the same face attribute, and Protocol-2, which examines generalization to novel face attributes. Six DeepFake detectors are chosen for comparison, including Lip-Forensics [28], SLADD [24], ICT [31], DCL [32], CFM [37], and SO-ViT-B [41]. According to Table IV, SLADD, DCL, and CFM perform poorly in the Protocol-1 test. This suggests that their learned features are manipulation-specific, thus falling short of detecting the same face attribute forgeries when altered by different methods. In stark contrast, semantics-oriented methods such as SO-ViT-B and the proposed SJEDD yield favorable results in the Protocol-1 test, underscoring the advantages of learning and integrating high-level semantic features at the global face attribute level with low-level signal features at the local face region level. When switching to the Protocol-2 test, SJEDD continues to outperform manipulation-oriented methods, particularly in generalizing to the unseen gender attribute. This success may be attributed to the learned transferable semantic and signal features from various local face regions and across different global face attributes, thereby facilitating generalization to novel face attributes.
TABLE IV: AUC results in the cross-attribute test setting. All models are trained on the training set of FF++ (and its augmented versions). “–” indicates that no such evaluation can be properly conducted
@@
SJEDD (Ours) 99.86 80.16 92.23 86.45 86.37 89.03 86.87
TABLE V: AUC results of SJEDD against six contemporary DeepFake detectors in the cross-dataset test setting. All models are trained on FFSC

### v2

@@ IV Proposed Method: SJEDD
In this section, we describe the proposed DeepFake detector SJEDD in detail, including 1) general problem formulation, 2) joint embedding for DeepFake detection, and 3) bi-level optimization for parameter estimation. The system diagram of SJEDD is shown in Fig. 3.
@@ IV-A Problem Formulation
Given an input face image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$, we assign to it a vector of binary labels $\bm{y}=\{0,1\}^{C}$, corresponding to one primary and $C-1$ auxiliary tasks. Specifically, $y_{1}=1$ signifies a forged face, and $y_{1}=0$ means the opposite. For $i>1$, $y_{i}=1$ indicates that the $i$-th face attribute/region is fake, and $y_{i}=0$ otherwise. Next, we define an unnormalized probability distribution over the label hierarchy:
@@
where $s_{i}(\bm{x})$ represents the raw score of the detector for the $i$-th label, and $\mathbb{I}[\cdot]$ is an indicator function. Eq. (1) emphasizes two label relations: 1) if the $j$-th node is activated (i.e., $y_{j}=1$), at least one of its parent nodes, collectively indexed by $\mathcal{P}_{j}$, must also be activated; and 2) if the $i$-th node is activated, at least one of its child nodes, indexed by $\mathcal{C}_{i}$, must also be activated. We then normalize Eq. (1) to obtain the joint probability:
@@ IV-B Joint Embedding
We now describe the text templates to encode the label hierarchy and implement joint embedding using vision-language correspondence to estimate the raw scores for all labels.
@@ IV-B1 Text Templates
We create the text templates based on face attributes, explicitly incorporating semantic face priors into the joint embedding space. For the root node face, the text template is defined as “a photo of a fake face.” At the global face attribute level, we use “A photo of a face with the global attribute of $\{g\}$ altered,” where $g$ belongs to the set of global face attributes {expression, identity, physical_inconsistency}. At the local face region level, we adopt “A photo of a face with the local $\{l\}$ region altered,” where $l$ is in the set of local face regions {eye, eyebrow, lip, mouth, nose, skin}. In total, we have ten candidate text templates.
@@ IV-B2 Vision-Language Correspondence
We construct two encoders: an image encoder $\bm{f}_{\bm{\phi}}:\mathbb{R}^{H\times W\times 3}\mapsto\mathbb{R}^{N\times 1}$, parameterized by $\bm{\phi}$, which computes the visual embedding $\bm{f}_{\bm{\phi}}(\bm{x})$ from an input face image $\bm{x}$, and a text encoder $\bm{g}_{\bm{\varphi}}:\mathcal{T}\mapsto\mathbb{R}^{N\times 1}$, parameterized by $\bm{\varphi}$, which produces a set of textual embeddings $\{\bm{g}_{\bm{\varphi}}(\bm{t}_{i})\}_{i=1}^{C}$. Here, $\bm{t}_{i}$ denotes the $i$-th text template, and $\mathcal{T}$ encompasses all candidate text templates. We estimate the raw score in Eq. (1) by
@@
The joint embedding approach eliminates the need for manual splitting of task-agnostic and task-specific layers. Through end-to-end optimization, it learns to allocate model capacity for each task automatically.
@@ IV-C Bi-level Optimization
Our goal is to optimize the joint embedding function, parameterized by $\bm{\phi}$ and $\bm{\varphi}$, which outputs the raw scores for all labels, based on which their marginal probabilities $\hat{p}(y_{i}|\bm{x};\bm{\phi},\bm{\varphi})$ can be computed. Given a training minibatch $\mathcal{B}_{\mathrm{tr}}=\{\bm{x}^{(k)},\bm{y}^{(k)},\mathcal{Z}^{(k)}\}_{k=1}^{K}$, where $\bm{y}^{(k)}=\{0,1\}^{C}$ is the complete ground-truth label vector and $\mathcal{Z}^{(k)}\subseteq\{1,\ldots,C\}$ is the index set of observed labels, we minimize a linear weighted sum of losses for each observed label:
@@
where $\bm{\lambda}=[\lambda_{1},\ldots,\lambda_{C}]^{\intercal}$ is the loss weighting vector to be optimized along with the model parameters $\bm{\phi}$ and $\bm{\varphi}$.
While the de facto cross-entropy loss can be used to implement $\ell(\cdot)$ in Eq. (5) for parameter estimation, it is unbounded from above [49]. This may result in excessive penalties for challenging training examples, thus hindering model capacity allocation and loss weighting adjustment in multitask learning. To address this, we use the fidelity loss [49]:
@@
To automatically adjust the loss weighting vector $\bm{\lambda}$ and to prioritize the primary DeepFake detection task, we employ bi-level optimization [48]. At the outer level, we minimize the primary task objective with respect to $\bm{\lambda}$ on the validation minibatch $\mathcal{B}_{\mathrm{val}}$. At the inner level, we minimize the linearly weighted loss with respect to $\bm{\phi}$ and $\bm{\varphi}$ on the training minibatch $\mathcal{B}_{\mathrm{tr}}$, giving rise to
@@
where ||$\cdot$|| ⋅ | denotes the cardinality of a set. As suggested in [44], we share loss weightings for tasks within the same level. That is, we learn to adjust three $\lambda_{i}$’s: one for the primary task, one for tasks at the global face attribute level, and one for tasks at the local face region level. We sample training and validation data as different minibatches in the same training dataset. We utilize the finite difference method as an efficient approximator to solve Problem (7).
TABLE II: AUC results of the proposed SJEDD against existing DeepFake detectors in the cross-dataset setting. All models are trained using the (respectively augmented) training set of FF++. We retrain CNND, Face X-ray, MADD, FRDM, SLADD, DCL, and SO-ViT-B to match the performance reported in their original papers. For SFDG, VLFFD, CCDFM, and Zhang24, we directly copy the results from their original papers due to the lack of publicly available implementations. The last column shows the mean AUC numbers, including/excluding the FF++ test set. The best results are highlighted in bold
@@
V Experiments
In this section, we first describe the experimental setups and then compare the proposed SJEDD with contemporary DeepFake detectors. In addition, we perform a series of ablation studies to validate the design choices of SJEDD from both the model and data perspectives.
V-A Experimental Setups
V-A1 Datasets
We adopt the widely used FF+ [20] and newly established FFSC [44] for training. FF++ contains $1,000$ real videos, manipulated by four DeepFake methods: Deepfakes (DF) [1], Face2Face (F2F) [46], FaceSwap (FS) [45], and NeuralTexures (NT) [4]. These videos come with three compression levels, and our experiments are conducted at the slight compression level, as recommended by [25, 74, 29]. We expand FF++ using the proposed dataset expansion technique described in Sec. III. In contrast, FFSC [44] organizes twelve face manipulations into a semantic hierarchical graph with five global face attribute nodes and six local face region nodes, requiring no further dataset expansion. We test DeepFake detectors on four independent datasets: CDF [51], FaceShifter (FSh) [50], DeeperForensics-1.0 (DF-1.0) [52], and DeepFake Detection Challenge (DFDC) [53].
V-A2 Implementation Details
SJEDD relies on the CLIP model [47] for joint embedding, where the image encoder is ViT-B/32 [100] and the text encoder is GPT-2 [101]. SJEDD is fine-tuned by minimizing Problem (7) using AdamW [102] with a decoupled weight decay of $10^{-3}$ and a minibatch size of $32$. The initial learning rate is set to $6\times 10^{-7}$, following a cosine annealing schedule [103]. The loss weightings are all initialized to one and optimized using Adam [104] with a learning rate of $10^{-3}$. Throughout training, the temperature parameter $\tau$ in Eq. (4) is fixed at $10$. Following [26, 21, 28], we randomly introduce one of five real perturbations to each input image with a probability of $0.3$: patch substitution, additive white Gaussian noise, Gaussian blurring, pixelation, and JPEG compression. These perturbations are constrained to preserve face semantics (i.e., label-preserving). The training process is run for up to $20$ epochs, taking approximately $21$ hours to complete on a single NVIDIA RTX 3090 GPU.
@@ V-A3 Competing Methods
We compare SJEDD against state-of-the-art DeepFake detectors: CNND [21], Face X-ray [23], MADD [28], Lip-Forensics [29], FRDM [30], RECCE [24], SBI [31], ICT [32], SLADD [25], DCL [33], CADDM [26], SFDG [34], VLFFD [35], CCDFM [98], MRL [36], CFM [38], Zhang24 [99], and SO-ViT-B [44]. CNND serves as a common baseline. MADD employs attention mechanisms for feature selection and aggregation. FRDM and CCDFM expose face forgery through high-frequency analysis. DCL, CFM, and Zhang24 leverage contrastive learning, while SFDG and MRL utilize graph learning. Face X-ray, SLADD, and CADDM incorporate manipulation localization as a manipulation-oriented auxiliary task. SLADD and VLFFD additionally include $C$-way manipulation classification. RECCE suggests reconstructing face images to learn generative face features. Face X-ray, SBI, ICT, SLADD, and VLFFD exploit different data augmentations, including face swapping between real faces [23, 31, 32] and local face region manipulation [25, 35]. Lip-Forensics and ICT rely on high-level lipreading and face identity features, respectively. Like the proposed SJEDD, SO-ViT-B employs a semantics-oriented multitask learning paradigm.
@@ V-A4 Evaluation Metrics
In line with [29, 30, 31, 32, 25, 26, 24, 33, 28, 27], we evaluate detection performance using the area under the receiver operating characteristic curve (AUC (%)).
@@ V-B1 Training on FF++
Cross-dataset evaluation. Table II shows the AUC results with several interesting observations. First, many detectors struggle to generalize. In contrast, SJEDD performs well on all five DeepFake datasets, particularly excelling in challenging datasets such as CDF, DFDC, and FFSC. Second, while SBI reports a high AUC of $93.18\%$ on CDF, it performs poorly on DF-1.0 and DFDC. This performance discrepancy may be due to the heavy reliance on manipulation-specific features, such as statistical inconsistencies in color appearances and landmark locations. Third, all competing methods score relatively low on DFDC, likely resulting from the significant domain shift caused by different imaging conditions and manipulation techniques. Last, SJEDD outperforms its predecessor, SO-ViT-B, confirming its improvements in model design (i.e., from discriminative learning to joint embedding) and loss function (i.e., from the cross-entropy to fidelity loss in bi-level optimization).
TABLE III: AUC results in the cross-manipulation setting. DF, F2F, FS, and NT represent Deepfakes [1], Face2Face [46], FaceSwap [45], and NeuralTexures [4], respectively. Gray numbers represent within-manipulation results
@@
SJEDD (Ours) 99.88 95.14 97.31 98.51 97.44
Cross-manipulation evaluation. We perform the cross-manipulation evaluation of SJEDD against six representative DeepFake detectors: MADD [28], RECCE [24], DCL [33], VLFFD [35], CFM [38], and SO-ViT-B [44]. Specifically, we train on one manipulation type from FF++ [20] (including additional augmented data) and test on the remaining manipulations from FF++. As shown in Table III, SJEDD surpasses other methods by clear margins, which we believe arises from two primary reasons. First, our semantics-oriented formulation helps SJEDD learn generalizable features across different manipulations that alter the same face attributes (e.g., NT $\rightarrow$ F2F, both modifying the global expression attribute and the local mouth and lip regions). Second, SJEDD models the joint probability distribution of face attributes (see Eq. (1)), thus promoting generalization across different face attributes (e.g., DF: identity $\rightarrow$ NT: expression). Once again, SJEDD outperforms its predecessor, SO-ViT-B, highlighting the effectiveness of the joint embedding formulation and the fidelity loss in bi-level optimization.
Cross-attribute evaluation. As suggested in FFSC [44], we adopt two semantics-oriented testing protocols: Protocol-1, which focuses on generalization to novel manipulations for the same face attribute; and Protocol-2, which examines generalization to novel face attributes. Six DeepFake detectors are chosen for comparison, including Lip-Forensics [29], SLADD [25], ICT [32], DCL [33], CFM [38], and SO-ViT-B [44]. The results on the FFSC main test set are shown in Table IV, where we find that SLADD, DCL, and CFM perform poorly under Protocol-1. This provides a strong indication that the features they learn are manipulation-specific and thus ineffective at detecting the same face attribute forgery by different methods. In stark contrast, semantics-oriented methods such as SO-ViT-B and the proposed SJEDD yield favorable results in the Protocol-1 test, underscoring the advantages of learning and integrating semantic features at the global face attribute level with signal features at the local face region level. When switching to Protocol-2, SJEDD continues to outperform manipulation-oriented methods, particularly in generalizing to the unseen gender attribute.
TABLE IV: AUC results in the cross-attribute setting using the FFSC main test set. All models are trained on the (respectively augmented) training set of FF++
@@
SJEDD (Ours) 99.86 80.16 92.23 86.45 86.37 89.03 86.85
TABLE V: AUC results of SJEDD against six contemporary DeepFake detectors in the cross-dataset setting. All models are trained on FFSC